OpenStudy (anonymous):

Need help with a Math test Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 6) and (1, 5) B. (10, 6) and (5, 1) C. (10, 4) and (5, 1) D. (10, 5) and (5, 1)

OpenStudy (greatlife44):

y = x-4 so technically when we plug in our x value we must get a y value that's the same as those ordered points.

OpenStudy (anonymous):

so b?

OpenStudy (greatlife44):

let's plug in x = 5 y = 1 y = 5-4 = 1 true can you see why?

OpenStudy (anonymous):

@staersiua what you understand by ordered pairs.

OpenStudy (anonymous):

yay i got it right

OpenStudy (anonymous):

thnx for all your help

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

nvm

OpenStudy (greatlife44):

@staersiua no problem, welcome to open study

rebeccaxhawaii (rebeccaxhawaii):

plug them in